0.800 watts of power are dissipated in a cell of emf 10.0 volts. If the internal resistance is 2.40 ȍ, find the resistance of the load and the current in the load. 


1
Expert's answer
2021-12-27T14:07:05-0500

"P=I^2R"


"0.8=I^2\u00d72.4"

"I=0.57735"


Lost voltage "=0.57735\u00d72.4"

"=1.38564V"


Terminal voltage "=10-1.38564"

"=8.61436"



Resistance of load "=\frac{8.61436}{0.57735}"

"=14.92052\Omega"



"\therefore" current in load and resistance of load

"=0.57735A"

and "14.92052\Omega"